1. Add the following numbers in the indicated bases (do not make any conversion to a different base):

a. 4607.548 + 1705.638

b. A6516 + F3216

c. 1100102 + 0101012

d. 4325 + 12345

1
Expert's answer
2022-05-03T03:31:52-0400

We have to add numbers bit by bit. If the sum in the row exceeds the base of the system, the value of this bit is equal to the base of the system minus the sum in the bit. 1 is added to the sum in the next digit.


a. 4607.54+1705.63=6515.37


b. A65+F32=1997


c.110010+010101=1000111


d.432+1234=2221